OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 8 cm. What is the approximate radius of the cylinder? A. 2.5 cm B. 4.0 cm C. 4.5 cm D. 20.4 cm

OpenStudy (anonymous):

@hartnn @ganeshie8

OpenStudy (anonymous):

pi*radius^2=64 radius^2=64/pi radius=sqrt(64/pi).

OpenStudy (anonymous):

Solve that.

OpenStudy (anonymous):

@eatacat I don't get it.

OpenStudy (anonymous):

Which part don't you get?

OpenStudy (anonymous):

Wait . . . square root of 20.3821656051 is 4.51466118387, so it must be C. I just didn't understand the formulas for a second. Thanks!

OpenStudy (anonymous):

It's C.

OpenStudy (anonymous):

I should've typed it out like this. r = √(64/pi)